OpenStudy (anonymous):

find parametric equations for the following: line y=3x-5 such that t=0 corresponds to the point (1,-2) on the line.

OpenStudy (dumbcow):

for linear functions, the parametric equations are usually x = t +C y = m(t + C) +b since t=0 refers to x=1 ..... C=1 x = t+1 y = 3(t+1) -5 = 3t -2